If two planes leave the same airport at 1:00pm, how many miles apart will they be at 3:00pm if one travels directly north at 150mph and the other travels directly west at 200mph?
Answer by checkley77(12569)

(150*3)^2+(200*3)^2=d^2
450^2+600^2=d^2
202,500+360,000=d^2
562,500=d^2
d=sqrt562,500
d=750 miles. ans. for the distance between these 2 planes @ 3:00 PM.